OpenStudy (anonymous):

HELP ALG 2 SERIES Find the 10th partial sum of the summation of 5i plus 2, from i equals 3 to infinity. 222 240 255 -355

OpenStudy (anonymous):

@yomylola

OpenStudy (anonymous):

You want \[\large\sum_{i=3}^{12}(5i+2)\] since \(i=12\) gives the tenth term in the series.

OpenStudy (anonymous):

These formulas should help: \[\large\sum_{i=1}^ni=\frac{n(n+1)}{2}\] \[\large\sum_{i=1}^n1=n\]

OpenStudy (anonymous):

I=3